### v1

# When Should We Introduce Safety Interventions During Pretraining?

Abstract
Ensuring the safety of language models in high-stakes settings remains a pressing challenge, as aligned behaviors are often brittle and easily undone by adversarial pressure or downstream finetuning. Prior work has shown that interventions applied during pretraining, such as rephrasing harmful content, can substantially improve the safety of the resulting models. In this paper, we study the fundamental question: “When during pretraining should safety interventions be introduced?” We keep the underlying data fixed and vary only the choice of a safety curriculum: the timing of these interventions, i.e., after 0%, 20%, or 60% of the pretraining token budget. We find that introducing interventions earlier generally yields more robust models with no increase in overrefusal rates, with the clearest benefits appearing after downstream, benign finetuning. We also see clear benefits in the steerability of models towards safer generations. Finally, we observe that earlier interventions reshape internal representations: linear probes more cleanly separate safe vs harmful examples. Overall, these results argue for incorporating safety signals early in pretraining, producing models that are more robust to downstream finetuning and jailbreaking, and more reliable under both standard and safety-aware inference procedures.
Language Models, Safety, Pretraining
1 Introduction
Large language models (LLMs) are increasingly deployed in high-stakes settings, such as medical domains (Thirunavukarasu et al., 2023; Liévin et al., 2024), social sciences and simulations (Ziems et al., 2024; Byun et al., 2025), and in robotics (Firoozi et al., 2025; Robey et al., 2024), where unsafe generations can cause real-world harms. Most production systems still rely on post-hoc alignment (e.g., RLHF (Ouyang et al., 2022), DPO (Rafailov et al., 2023), or constitutional variants (Bai et al., 2022)) that reshape surface model behavior after completing pretraining. Yet, a growing body of evidence suggests these methods are brittle: alignment effects are often restricted to the first few tokens of a response (Qi et al., 2024b) and are easily circumvented by benign finetuning (Qi et al., 2024a) or adversarial pressure through prompting (Zou et al., 2023; Chao et al., 2023). Misaligned behavior can be further elicited, and even amplified, by malicious or adversarial finetuning on a single task, which has been shown to generalize to other harmful tasks (Betley et al., 2025).
Figure 1: Attack success rate of the resulting base models as we vary the time at which we introduce safety pretraining interventions. For standard inference (Top-K), we report the shaded region as the standard deviation computed over 5 seeds. When using SafeBeam, the results are deterministic. Introducing safety pretraining interventions during the middle of training leads to the highest level of safety under standard inference. SafeBeam benefits from interventions being introduced at the start of pretraining.
As such, recent work has explored removing all harmful content from pretraining data via filtering approaches (O’Brien et al., 2025). Other work, however, has demonstrated that this could lead to a lack of understanding of harmful content, making models more prone to comply with harmful requests (Li et al., 2025). One recent work of interest (Maini et al., 2025) has explored various synthetic pretraining interventions to (i) maintain information and content from harmful pretraining examples, while (ii) removing explicitly harmful content or providing sufficient context to understand why particular pretraining data may be harmful. Safety pretraining interventions show promise for safer language model behavior, in terms of refusal to comply with harmful queries, adversarial jailbreaks, and even after benign finetuning.
Despite this progress, a central question remains underexplored: when during pretraining should safety interventions be introduced? Curriculum learning (Bengio et al., 2009; Graves et al., 2017) suggests that the order of examples (in terms of topic or difficulty) can qualitatively change downstream performance. However, such topics have been demonstrated primarily in non-language domains, with some work studying the impact of data mixture components at a more coarse-grained level (Xie et al., 2023; Jiang et al., 2024; Ye et al., 2025) or other work introducing benefits of different stages during pretraining (Blakeney et al., 2024). Intuitively, safety interventions introduced only at the end may be too superficial and be easily washed away by downstream finetuning applications, while signals introduced too early or too strongly may degrade standard capabilities and helpfulness or may over-regularize the model.
In this paper, we empirically study the impacts of introducing safety pretraining data at different times during pretraining. We adopt the various pretraining interventions from the work of Maini et al. (2025), such as synthetic rephrasing and contextualization, metadata annotation, and refusal training, and introduce them at different stages (i.e., after 0%, 20%, or 60% of the total amount of 600B tokens) of pretraining. All of our experiments revolve around the experimental setting of pretraining a 1.7B parameter language model with a SmolLM2 architecture (Allal et al., 2025) on a pretraining dataset containing FineWeb-Edu (Penedo et al., 2024), math, code, and other synthetic data (Allal et al., 2025). To evaluate the impact of timing of these interventions, we analyze the safety of the resulting models along multiple axes (i.e., base models, instruction-tuned models, and models after finetuning) and analyze how well the models separate harmful and safe content in their representation space.
Figure 2: Attack success rate of the resulting benign finetuned models on math data (GSM8k) as we vary the time at which we introduce safety pretraining interventions. For standard inference (Top-K), we report the shaded region as the standard deviation computed over 5 seeds. When using SafeBeam, the results are deterministic. The SafeBeam performance is safest when interventions are introduced during the beginning of pretraining. Standard inference benefits from introducing safety interventions early (i.e., 0% or 20%) during pretraining.
We generally find that introducing such interventions earlier leads to greater robustness of base models (i.e., before any instruction tuning) and to greater robustness of instruction-tuned models after a round of benign finetuning. Particularly, we see larger impacts when using a modified inference-time algorithm for safer generations (i.e., a safety-focused beam search). This is complemented by our representation analysis, where we explore the impact of the timing of such pretraining data on the notion of safe versus unsafe data in representation space. In those results, we also find that introducing such topics earlier leads to a greater distinction between harmful and safe content in the representation space, translating to a linear probe yielding higher performance at separating these classes in the last layer representation space of the trained language models. Finally, we find that introducing safety interventions early (i.e., at 20% or 60%) during pretraining leads to significantly more robustness under a threat model of an adversarial jailbreak learned via GCG (Zou et al., 2023), when compared to no safety interventions or safety interventions introduced at the beginning of training (i.e., 0%). Overall, our work suggests the incorporation of safety pretraining interventions earlier during training, yielding improvements in robustness to jailbreaks and to harmful user queries after benign finetuning.
2 Related Work
Pretraining Safety
Most language model pretraining employs simple heuristic filters that remove a small amount of harmful content, such as from particular URLs or explicit profanity (Raffel et al., 2020; Xue, 2020). One early work studied the role of human preferences in language model pretraining (Korbak et al., 2023), finding that it yields lower toxicity scores relative to finetuned models (on the 100M parameter scale). Other works have taken a more explicit data curation approach. More recently, works have studied heuristic filtering based on educational content (Gunasekar et al., 2023; Penedo et al., 2024), which also correlates with (lower levels of) toxicity. Yet, Vidgen et al. (2024) note that harmful examples remain even in these educational datasets. As such, some work explores the impacts of filtering unsafe content from the pretraining corpus, training only on safe examples (Li et al., 2025; O’Brien et al., 2025).
Most recently and relatedly, the work of Maini et al. (2025) studies the role of synthetic data generation to improve levels of safety in pretraining examples, while still retaining as much information as possible, finding large improvements in robustness after a round of benign finetuning. We expand upon this work to study the question of when such interventions should be deployed during pretraining.
Figure 3: Attack success rate of the resulting instruction-tuned models as we vary the time at which we introduce safety pretraining interventions. For standard inference (Top-K), we report the shaded region as the standard deviation computed over 5 seeds. When using SafeBeam, the results are deterministic. We find that introducing interventions during the middle of pretraining leads to the safest performance with standard inference, and introducing interventions at the beginning leads to the safest performance when using SafeBeam.
Curriculum Learning
Curriculum learning studies how the order and composition of training data shape generalization. Early work argues for presenting examples in a progression (e.g., from simpler to harder), and for adapting the sampling policy as training proceeds (Bengio et al., 2009; Graves et al., 2017). Such approaches have led to success in various settings, such as question-answering (Sachan and Xing, 2016) or reinforcement learning (Zhang et al., 2020). Other work has explored when and what types of certain curriculum learning strategies work, demonstrated on small-scale image classification tasks (Wu et al., 2021). Progress in the language model pretraining realm has been relatively static, likely due to the difficulty in measuring the hardness of content (Abdin et al., 2024), and perhaps even similarity between content (Sam et al., 2025a), as well as the computational demands of pretraining.
Stages of Language Model Pretraining
In large-scale language pretraining, most practical curricula have focused on adjusting mixture weights over data sources rather than token-level difficulty, largely due to computational complexity. For example, optimizing source mixtures during pretraining can yield gains in downstream performance and stability (Xie et al., 2023), and a line of work explores adaptive sampling or stage-specific mixtures that emphasize more useful distributions as training evolves (Jiang et al., 2024; Ye et al., 2025). Complementary results in a more coarse-grained setting suggest that pretraining in distinct stages can change what models retain and how they transfer (Blakeney et al., 2024). Our study is aligned with this perspective but targets a safety-specific question: instead of asking which domains to upweight, we vary when safety-augmented data enters pretraining. This isolates timing effects and lets us measure whether safety signals introduced earlier versus later have a greater impact.
3 Curricula for Safety
This section formalizes our notion of a safety curriculum, detailing the particular interventions and how/when they are incorporated into the pretraining process.
3.1 Safety Pretraining Interventions
Crucial to this work are safety interventions that can be explored during the pretraining process, which are introduced in the work of Maini et al. (2025). We briefly introduce and describe these various pretraining interventions below.
Contextualized Rephrasing
For pretraining examples that contain harmful content or instructions, the example is rewritten to preserve useful factual or domain knowledge while explicitly providing context that frames the material as risky or inappropriate to execute. The rewritten passages describe why the behavior is unsafe, when it is permissible to discuss at a high level, and how to redirect toward safer generations. The goal is to retain conceptual coverage (so the model “knows about” the topic) while discouraging unsafe execution, all within the same pretraining objective. Such examples are rephrased with a LLaMA3.1 8B parameter model (Touvron et al., 2023).
Refusal Training
The most harmful pretraining examples are converted with a similar synthetic data pipeline into conversation-style request and refusal pairs to incorporate throughout pretraining. Prompts that elicit harmful actions are paired with refusals that model helpful-but-safe behavior: declining the request, briefly explaining safety rationales, and, when appropriate, offering a safe alternative or high-level guidance. These pairs are added directly to the pretraining corpus so that refusal behavior is learned natively to the model throughout pretraining, rather than only during a post-hoc alignment phase. This strengthens robustness towards downstream finetuning, where refusal behavior can easily be eroded either on benign or explicitly harmful tasks (Qi et al., 2024a; Betley et al., 2025).
Figure 4: Comparison of overrefusal rates on Alpaca (Taori et al., 2023) when using standard (Top-k) sampling for inference. We find that incorporating interventions earlier during pretraining leads to a better compliance rate on benign requests.
Metadata-annotated Pretraining and SafeBeam
Examples are flagged with a special token (e.g., “<potentially_unsafe_content>”), which indicates that a pretraining example is harmful. As with the other interventions, this is implemented as ordinary text (randomly inserted with a 5% chance at each of the token positions of the pretraining example) that the model predicts, avoiding changes to the training loss. These metadata annotations build associations within the model of harmful content and the special token, which enables steerability during inference-time. Specifically, this enables a SafeBeam search algorithm, which is a safety-focused inference-time modification. This modifies a standard beam search algorithm by dropping candidate beams that have a high probability of the next token being the harmful special token, leading to safer generations. The ultimate goal of such an intervention is to steer the model towards safer and high-quality completions (Yuan et al., 2025).
We adopt the same resulting datasets from applying these interventions on the FineWeb dataset (Penedo et al., 2024) and other components of the SmolLM2 pretraining corpus (Allal et al., 2025).
3.2 Safety Curricula
A central question of this work is how the timing of safety interventions during pretraining shapes downstream robustness. Prior work on safety pretraining (Maini et al., 2025) has introduced interventions from the very start of training—effectively modifying the full corpus before pretraining begins. Other work has demonstrated that orderings of certain concepts can improve the speed of learning (Chen et al., 2023). Similarly, in the human learning process, we often first learn about general knowledge, such as language and arithmetic, before addressing potentially harmful or controversial content, such as wars and tragedies. At the other extreme, filtering approaches (O’Brien et al., 2025) that permanently and completely remove all harmful content during pretraining potentially lose too many capabilities from the underlying model. We adopt a similar approach here, where we study a curriculum of first training on safe-only data before introducing safety interventions after a certain percentage of tokens of the full 600B token pretraining budget.
To study the role of the timing of introducing safety interventions, we study 4 main settings. The first setting (0%) applies interventions from the very beginning, which directly mirrors the setup in prior Safety Pretraining work (Maini et al., 2025). This captures no pretraining phase on safe-only data. The second setting (20%) delays interventions until after the model has already seen a substantial portion of “safe” general web-scale data, allowing it to acquire broad linguistic and factual knowledge before being guided by safety signals. The third setting (60%) pushes this even further, introducing interventions only in the later stages of training, when much of the model’s base distribution of language has already been established. Finally, the last setting (100%) corresponds to training completely on safe-only data, which is effectively equivalent to filtering unsafe content for the entirety of pretraining. This has been explored in prior work (Maini et al., 2025), as well as in more harm-specific settings such as biological risks (O’Brien et al., 2025).
To also vary the types of interventions employed, we explore the two primary settings from Safety Pretraining (Maini et al., 2025): (i) we first study using all of the interventions in conjunction, including the metadata tagging and inference-time modifications, (ii) using only rephrasing and refusal training without any metadata tagging. The second set of experiments here considers a slightly less aggressive modification to pretraining, only intervening in terms of data curation.
Figure 5: Comparison of overrefusal rates on Alpaca when using the SafeBeam inference algorithm. We find that introducing pretraining interventions and metadata earlier during pretraining leads to more helpful compliant behavior, with a slight increase in overrefusal rates.
4 Experiments
To evaluate the impacts of the timing of safety curricula, we adopt evaluations along multiple axes of the model pipeline. Our evaluations include safety (of base model, instruction-tuned models, and finetuned models), helpfulness on harmless and benign queries, jailbreakability through adversarial suffixes, and in analyzing model internals and the underlying representation space.
4.1 Safety Evaluations
We follow the work of Maini et al. (2025), performing safety evaluations at 3 main stages: (i) base model – immediately after pretraining, (ii) instruction-tuned models – after a standardized post-training setup, and (iii) benign finetuned models – after a round of benign finetuning on math data, e.g., GSM8K (Cobbe et al., 2021).
We first evaluate the safety of our models immediately after pretraining. We use the completion-style harmful request benchmark from Maini et al. (2025). This evaluates the tendencies of the base models to complete the beginning of a harmful prompt. Next, we evaluate the models after performing a standard instruction tuning recipe, which also includes some safety training and refusal examples. Our evaluation is comprised of 4 standard safety benchmarks: HarmBench (Mazeika et al., 2024), TDC (Maloyan et al., 2024), JailbreakBench (Chao et al., 2024), AdvBench (Zou et al., 2023). We also incorporate a harder evaluation setting that incorporates adversarial jailbreaking, where we learn adversarial suffixes via the GCG algorithm (Zou et al., 2023) that are optimized on each model to increase the rate of compliance with harmful requests.
Finally, we evaluate the models after an epoch of benign, supervised finetuning on a helpful, non-adversarial dataset of GSM8K (Cobbe et al., 2021), which is common practice for mathematical reasoning, without any focus on safety or alignment. Despite this being non-adversarial data, other works have demonstrated that such rounds of supervised finetuning can erode aligned behaviors (Qi et al., 2024a; Betley et al., 2025).
4.2 Helpfulness Evaluations
Whenever introducing safety interventions (that lead to potentially increased refusal rates), it is crucial to also assess any impacts on the helpfulness of the resulting models. We evaluate the performance of our models on complying with benign requests from the Alpaca dataset (Taori et al., 2023), using GPT-4o-mini as a judge to evaluate if responses are (1) compliant and helpful, (2) compliant and unhelpful, (3) partial refusal, or (4) full refusal. Desirable responses are categories (1) and (2), representing model compliance with varying levels of helpfulness. Categories (3) and (4) denote overrefusal in this setting.
When evaluating models that use the modified SafeBeam inference algorithm, we also add in a category (5) shortcircuit, which stops generation when the harmful metadata token is produced. This category reflects a separate and practical setting where we can post-process model responses that contain the harmful token at production time, replacing them with a default refusal message taken from a set of known safe completions.
Figure 6: Jailbreakability of our models via the GCG algorithm (Zou et al., 2023) as we vary the timing of safety pretraining interventions. We find that introducing interventions during the middle stages of pretraining greatly improves the steerability of models towards safe completions and refusals when exposed to adversarially learned suffixes.
4.3 Representation Analysis
To analyze the impact of various curricula on the underlying representation space of our models, we evaluate the separability of safe vs harmful content. We train linear probes to classify between examples from FineWeb data that have been annotated as safe versus unsafe, learned over the representations taken from the internal activations of the model’s final layer. For the safety label, we use the safety score from Maini et al. (2025), which is produced by GPT-4o-mini that judges pretraining examples with a safety rubric.
4.4 Training Details
Pretraining Setup
For our experiments, we adopt the same setup of the SmolLM2 (Allal et al., 2025) for all our experiments. We train 1.7B parameter language models with a pretraining corpus of a mixture of FineWeb-Edu, StackOverflow, FineMath, and Cosmopedia. For our safety pretraining interventions, we adopt the modified datasets of SafeWeb, RefuseWeb, and MoralEducation from Maini et al. (2025), which are safety-focused, rephrased and recontextualized versions of harmful pretraining examples from FineWeb-Edu.
Pretraining is performed using LitGPT (AI, 2023), with FlashAttention-2 enabled and mixed-precision training for efficiency. We use the same hyperparameters (e.g., learning rate schedule, batch size, and sequence length) as in the SmolLM2 pretraining setup.
Post-training Setup
For a standardized post-training stack, we use a combination of Ultrachat-200k with AllenAI WildGuardMix and WildJailbreak datasets for safety instruction tuning. This follows standard practice from prior works in safety training (Zou et al., 2024; Maini et al., 2025).
5 Results
We now present our findings on how safety curricula influence robustness, helpfulness, and internal representations.
Figure 7: Separability of safe vs unsafe content (as annotated by GPT-4o-mini) in the last layer representation space of base models (Left) and instruction-tuned models (Right), as we vary the start timing when we incorporate safety pretraining interventions. We find that introducing interventions earlier yields greater separability in representation space (especially with the incorporation of harmful metadata tags). The shaded regions denote the standard deviation computed over 5 training runs.
5.1 Safety Results
Base Models
Introducing interventions earlier in pretraining consistently improves the safety of base models (Figure 1). We find that introducing safety pretraining interventions during the middle of training (60% or 20%) leads to the safest base models in terms of unsafe completion rate with standard top-k sampling. Not introducing any safety pretraining interventions and training on safe-only data (100%) leads to the most unsafe performance. Our results also suggest that an earlier inclusion of metadata (at 0%) leads to measurable gains when using SafeBeam inference.
Instruction-tuned Models
A similar pattern holds with the models that have undergone instruction tuning (Figure 3). Models trained with interventions during the middle of pretraining exhibit stronger safety performance than those with no interventions. Again, SafeBeam benefits the most when metadata is incorporated from the beginning (i.e., 0%).
Benign Finetuned Models
The clearest and most noticeable effects emerge in our evaluation setting after a benign finetune on mathematical data (Figure 2). Here, SafeBeam achieves the greatest robustness when interventions are included from the very start of pretraining. For standard top-k decoding, introducing interventions after roughly 20% of training yields the strongest post-finetuning safety.
Taken together, our evaluations suggest two main takeaways. First, when using modified SafeBeam inference, introducing metadata-annotated unsafe data earliest in pretraining leads to the best steerability towards safer completions. Secondly, with top-k inference, introducing safety pretraining interventions during the middle of training (either 20% or 60%) leads to the most robust performance. This suggests that allowing the model to first absorb general-purpose safe web-scale data before applying any safety interventions may produce more durable alignment when subsequent benign finetuning is performed.
5.2 Overrefusal Results
In assessing the impact of the pretraining interventions on the helpfulness of the models, we find that when using standard top-k sampling (Figure 4), there is a slight improvement when incorporating interventions earlier during pretraining (e.g., at 0% or 20%). This is likely due to models being able to learn more efficiently from high-quality synthetic data (Maini et al., 2024). Overall, this suggests that safety pretraining interventions observe no (or only minor) tradeoffs in overrefusal.
When using the SafeBeam algorithm (Figure 5), we find that incorporating pretraining interventions earlier leads to slightly higher overrefusal rates; however, within the compliant responses, helpfulness increases. Therefore, while overall refusals do increase, the quality of compliant responses improves, suggesting improved performance on general instruction-following tasks. A likely hypothesis for this behavior is that during SafeBeam, the most potentially harmful beams are discarded, and the remaining beams are selected to generate model responses. When pretraining interventions are included earlier during training, the model likely better learns the correspondence between the harmful metadata tag and unsafe content, as well as safer generations through the safe synthetic pretraining data. This leads to the remaining beams in SafeBeam being of higher quality and, ultimately, leading to more helpful responses.
5.3 Jailbreaking Results
We also present results on how susceptible our models are to jailbreaking attacks using the GCG algorithm (Zou et al., 2023), as we vary the timing of the safety pretraining interventions. Surprisingly, we find that introducing pretraining interventions after some initial safe-only pretraining (i.e., at 20% or at 60%) leads to significantly more robustness to jailbreaks via GCG when using the SafeBeam inference algorithm (Figure 6). On the other hand, we find that jailbreakability under standard inference does not noticeably change given different intervention timings.
5.4 Representation Results
In settings without harmful metadata tags, earlier interventions slightly improve the separability of safe versus unsafe points (with a linear classifier over the last-layer activations). When metadata is included, the effect is more pronounced: introducing interventions from the start yields the highest linear-probe AUC, suggesting a much cleaner separation of harmful and safe content in the representation space. This also supports the improved SafeBeam robustness, where models are more steerable towards safer responses when interventions are included early in pretraining.
6 Conclusion
Our work studies the role of different safety curricula, or how the timing of safety pretraining interventions (e.g., when and how safety-augmented data enters pretraining) impacts the robustness of our models. Our experiments show that incorporating safety pretraining interventions indeed help, and the clearest result is that there is much improved robustness after benign finetuning when pretraining interventions are introduced earlier (e.g., at 0% or 20% of the pretraining tokens). This also manifests into impacts on the model’s underlying representation geometry; incorporating interventions and metadata earlier in pretraining leads to greater separation of safe vs unsafe content. Taken together, these findings highlight how the timing of safety interventions is an important choice in shaping a model’s generations and internal representations, offering guidance for building more trustworthy and socially responsible language models.
Appendix A Additional Experiments
Figure 8: Separability of safe vs unsafe content (as annotated by GPT-4o-mini) in the first layer of base models (Left) and instruction-tuned models (Right), as we vary the start timing when we incorporate safety pretraining interventions. The shaded regions denote the standard deviation computed over 5 training runs.
Figure 9: Separability of safe vs unsafe content (as annotated by GPT-4o-mini) in the middle (13th) layer of base models (Left) and instruction-tuned models (Right), as we vary the start timing when we incorporate safety pretraining interventions. The shaded regions denote the standard deviation computed over 5 training runs.
Additional Separability Results
In the main text, we present results on separability of harmful vs safe content in the final layer of the resulting pretraining models. Here, we also present results for the first layer and the middle (13th) layer. We observe similar trends, although we find greater separability in the final layer when compared to earlier layers in the models.
Appendix B Additional Experimental Details
Evaluation
For our standard inference evaluations, we use top-k sampling with $k=10$ and truncate after 256 tokens. We average the results over 5 samples. For SafeBeam evaluation, we run a beam search with $10$ beams for 256 tokens, and discard the top 50% of unsafe beams (in terms of probability of the next token being the harmful metadata token). We follow the same exact setting as Maini et al. [2025]. We adopt the same GPT-4o-mini judge and judging prompt for helpfulness evaluations, and the same HarmBench judge and judge prompt [Mazeika et al., 2024] for our safety evaluations.
Adversarial Jailbreaking
For computing the adversarial suffixes via the GCG algorithm, we learn an adversarial suffix on the target model, optimizing for the target string of “Sure”. We perform optimization for 500 steps, with an initialization of “x x ... x”, and top_k of 256.
Finetuning
For our supervised instruction tuning, we finetune with a batch size of 4 with a learning rate of 2e-5 for one epoch, using cosine learning rate decay with a warmup ratio of 0.03. For our benign finetuning experiments, we perform supervised finetuning on GSM8k for a single epoch with a learning rate of 2e-5 with the same learning rate schedule as in instruction tuning.
Separability Experiments
In our separability experiments, we use the last token position and the final layer of all models. For each model, we train a logistic regression classifier on a balanced subset of 500 safe and 500 unsafe examples from the training split, and the trained probe is evaluated on the full balanced test set. We report the mean, standard deviation, and standard error, when computed over five seeds, of the test AUC (area under the ROC curve).
Appendix C Additional Discussion
One limitation of our experiments is that the introduction of safety pretraining interventions at different times during pretraining means that certain models have seen a larger number of tokens for synthetic contextualized data or refusal data. Thus, our results do not control the exact number of tokens from each pretraining mixture source. Finding an optimal “ordering” for a curriculum remains an interesting open question.
We also note that the notion of safety that we focus on in this paper is rather general, taken from the taxonomy proposed by the MLCommons [Vidgen et al., 2024]. This captures a wide variety of potential harms and undesirable behaviors. Our method likely requires additional efforts to tackle safety for specific capabilities, such as bio-weapons and cybersecurity [Li et al., 2024] or reliability of agents in context-dependent scenarios [Andriushchenko et al., 2025, Sam et al., 2025c].
We also note that our results on the separability of safe vs unsafe content has greater implications for the monitoring of such models. These experiments contribute to a line of work that develops external monitoring to predict when undesirable generations will occur and steer towards safer outputs [Turner et al., 2023, Zou et al., 2024]. This may even have implications for complementary alternative approaches that are adopted in black-box settings and use external models [Sharma et al., 2025, Sam et al., 2025b], which is an area for future study.
Appendix D Computational Resources
To perform pretraining for each of the 1.7B parameter models on 600B tokens, we used 4 nodes of 8xH100 GPUs for roughly 6-7 days. Finetuning requires roughly 4 H100 GPUs for less than 4 hours. Evaluations are run on a single NVIDIA L40S GPU in less than an hour.